a particle moving with uniform acceleration has its final velocity [v=(150+8x)1/2] m/s where x is the distance travelled by the body. Then the acceleration is 
​1) + 4 m/s2
​2) – 4 m/s2
​3) + 8 m/s2
​4) – 8 m/s2
Answer: + 4 m/s2

Grade:11

1 Answers

Eshan
5 years ago
v=\sqrt{150+8x} \implies v^2=150+8x

This can be compared to equation of motion for constant acceleration,v^2=u^2+2ax

Hencea=+4m/s^2